# Mastering Board Games by External and Internal Planning with Language Models

### v1

@@ 3.1 External Search with State-Tracking MAV
We adapt MCTS (see Algorithm 1) to use the learned MAV model to predict the state transition from the string description of a parent node $\mathcal{N}_{t-1}(s)$ and action $a_{t-1}$ to the child node $\mathcal{N}_{t}$, as shown in Figure 3. We use the obtained information to store a (predicted) string description $s_{t}$ of the child node, which is then expanded with legal actions $\boldsymbol{a}_{t}^{L}$ and their associated state-action values $\boldsymbol{Q}^{(i)}(s_{t},\boldsymbol{a}_{t}^{L})$ used to compute a prior according to Equation 1 and the child’s value as the maximum over all state-action values, i.e.$\max_{a_{t}\in\boldsymbol{a}_{t}^{L}}Q^{(i)}(s_{t},a_{t})$.
@@
The MAV model may hallucinate and return responses that are improperly-formatted. To address this, for each game, a parser function is responsible for translating the output into the next state, legal actions, and their values. If the parsing fails – which can only happen if a response violates a predefined format – then the evaluation at the node is marked as a hallucination and a special value of $\infty$- ∞ is assigned to the node to avoid future consideration. This procedure makes state-tracking MCTS robust to hallucinations, by explicitly avoiding states where the hallucination occurred.
3.2 Performance Considerations and Async MCTS
Figure 4: An overview of Asynchronous MCTS.
Since modern large transformer models can take orders of magnitude longer for a single inference than an optimized deep convolutional network (such as the ones used in MuZero (Schrittwieser et al., 2020)), keeping the elapsed time for an MCTS search to a reasonable level can be a challenge. The inference time of a single model call dominates the time take for a search, so we aim to keep the total calls low. Since the prior and value for a new node are both obtained for the MAV, there is at most one model call per simulation of MCTS. In the low simulation count regime that we operate in, reusing the search trees is a simple optimization that reduces the number of server calls by 5-10% in our experiments.
Despite this lower number of inference calls, it can still be costly to make many model calls. Hence, we run an asynchronous version of MCTS that is conceptually very similar to Algorithm 1, but with the main loop (between lines 79) changed. An overview is shown in Figure 4.
Parallel MCTS implementations usually rely on virtual losses (Mirsoleimani et al., 2017; Chaslot et al., 2008) to avoid multiple threads choosing the same simulation paths down due to deterministic action choices. Our asynchronous MCTS introduces three new parameters:
Batch size $b$.
The number of simulations, each leading to an evaluation (model call), that are done asynchronously at any given time.
Timeout $t_{0}$.
Virtual Count Value $n_{c}$.
The a “virtual” value to temporarily add to $N(s,a)$ during the simulation to discourage all simulations in a batch from choosing the same path.
In Asynchronous MCTS, simulations are instead performed in batches. First, $b$ downward passes are performed serially to queue $b$ evaluations. The $b$ evaluations are performed in parallel and the MAV outputs are returned asynchronously.
The evaluations are then processed one-by-one as they return. For each queued evaluation that was successful (i.e., did not time out), its values are then backpropagated up the tree (serially). The state-action pairs that are visited before the queuing of evaluations are assigned temporary virtual counts: their visit counts are temporary set to $N(s,a)=N(s,a)+n_{c}$. The visit counts and virtual counts are then subtracted upon the upward pass, i.e.$N(s,a)=N(s,a)-n_{c}$.
Dynamic Virtual Counts
@@ 3.2 Performance Considerations and Async MCTS
In our low simulation count settings, using the virtual losses accompanied by the fixed virtual counts $n_{c}$ did not strike a satisfactory balance between exploration and exploitation. To circumvent this obstacle, we introduce the dynamic virtual counts that dynamically assign more weight to the virtual count values closer to the leaf nodes. Suppose simulation $m$ encounters leaf node $\mathcal{N}_{l}$. We define a virtual count for states-action pairs $(s_{t},a_{t})$ visited in simulation $m$ and leaf $\mathcal{N}_{l}$
@@
Figure 5: Dynamic virtual counts for $n_{\text{max}}=8$ and $n_{\text{min}}=2$. Virtual counts from separate simulations are added, as is the case in the $+4$ node with two children ($+4=(+2)+(+2)$).
@@ 4 Internal Search
The prompt for internal search resembles that of the MAV models, but includes a preamble with search parameters (tree depth and breadth, see Figure 6). The format of the target data is inspired by depth-first order traversal of minimax trees, as an iterative and linearized sequence of minimax tree traversal. Hence, following this format corresponds to an algorithmic execution task, akin to CLRS-Text (Markeeva et al., 2024).
The training data is based on target search trees, which were constructed using depth 3 (N.B., depth-zero is MAV), by annotating states (e.g. chess states being annotated with Stockfish), and expanding the top 5 moves into trees. This results in high quality target search trees, similar to those internally generated by engines such as Leela or Stockfish. To diversify training data and enable search budget control for trained models, prompts were composed with diverse search parameters, ranging depth 1–3 and breadth 2–5, and continuations yield corresponding trees that were subsampled to match the parameters while fitting into context size. This necessitated excluding the biggest parameter combination (i.e., omitting depth 3, breadth 5 examples).
An example internal search trace predicted by the trained model, along with the corresponding search tree, are shown in Figure 6.
@@
We leveraged the pre-trained MAV model and fine-tuned it using a mixture of $60\%$ MAV data and $40\%$ search data. Fine-tuning was run for 20,000 steps, using a batch size of 512.
(a) External search results. We vary the number of simulations / MAV model calls.
(b) Internal search results. We vary the breadth ($b$) and depth ($d$) of the minimax search performed in a single model call. The $x$-axis shows the number of tokens in the prompt+response.
Figure 7: Chess Elo improves with more test-time compute.
@@ 5.1 Evaluation
Evaluating large language models in general and reasoning specifically is a broad field (Chang et al., 2024), within which games have been established as an evaluation benchmark (Costarelli et al., 2024; Duan et al., 2024). We evaluate language models in a games league which runs our agents in head-to-head match-ups, sampling combinations uniformly at random from a pool of agents. We report internal Elo ratings (relative Elo only between members of the population) as well as external Elo ratings in cases where some of our agents have externally reported Elos.
For chess evaluation, we rely on the state-of-the-art engine Stockfish at different playing strengths to estimate the widely used external Elo rating. This is done by calibrating the internal Elo with externally-reported ratings using a linear fit. As an additional, we include the Ext-BoN model (Ruoss et al., 2024), which showed Grandmaster-level performance on chess with Transformers (see Appendix for details). We use a set of TCEC (Top Chess Engine Championship) opening positions (tce, ) (Table LABEL:tab:long). In each match-up between two agents, a specific opening is used, and agents swap seats to ensure each agent plays each opening both as black and as white. Every instance of Stockfish is run with 2 seconds of search time. We delineate further our evaluation setting together with the closest related works in the Appendix.
@@
Table 2: Performance comparison of different agents in Chess, Chess960, Connect Four, and Hex. In MAV–MCTS agents, $M$ refers to the number of simulations. Important: The external Elos of the Stockfish agents are anchored to the CCLR Blitz Elos. Therefore, the Elo estimates for the MAV–MCTS agents are not accurate as those agents would not be able to play at blitz time controls.
@@
Table 3: Error rate analysis for MAV model in chess positions. We consider positions drawn from Lichess puzzles, and from a dataset of randomly generated boards that never occur in the course of normal gameplay, and are highly out-of-distribution compared to the training data.
@@ 5.2 Multi-Action-Value Results
In terms of chess playing strength, MAV reaches an external Elo of 2923 when using mean scoring and 2875 when using max scoring. It thus outperforms several strong baselines, including Stockfish–L10 and performs comparably with the Ext-BoN model by Ruoss et al. (2024). For Chess960 MAV outperforms Ruoss et al. (2024), perhaps due to being trained explicitly on Chess960 data.
In Table 3, we analyze the legal move rate, the precision and recall of the predicted “top-all” moves, and the accuracy of predicting the next FEN state after applying a chess move. The results demonstrate that MAV is able to reliably perform all of these actions.
@@
During the opening and endgame, human players often rely on memorized opening lines and endgame theory, while middlegame requires more calculation and intuition. We can see the same pattern in the games of MAV (see Figure 8). Overall, 10% of the positions played by MAV in evaluation appear in its training data, while between moves 20 and 50, virtually no position has been seen by MAV during training. These results show that in order to avoid losing games during middlegame, MAV is required to generalize.
[…]
@@ 5.3 External search results
For our external-MCTS agents, we tune hyper-parameters using a combination of manual and head-to-head comparisons. We found that setting $\tau=1.0$, $k=5$, and $\varepsilon=0.1$ worked best for the prior. Please see the appendix for details. For Async MCTS we used a batch size $b=16$. We found that lower batch sizes were always preferred in terms of performance. After hyper-parameter tuning, we used dynamic virtual count values $n_{\text{min}}=2$ in all cases, and $n_{\text{max}}$ varying with the number of simulations: $(M,n_{\text{max}})\in$$(100,8)$, $(250,8)$, $(500,8)$, $(1000,16)$, (2000,32)}(2000,32)\}( 2000 , 32 ) }. We also include a basic MCTS baseline, which refers to MCTS with a uniform prior and random rollouts run with 100 simulations.
We first analyze the performance of the scoring method for the value function when used within MAV–MCTS . We do this by running a large tournament amount all MAV–MCTS variants, various levels of Stockfish, MAV on its own, and basic MCTS. MAV–MCTS with mean scoring performs better in all cases, with a 25-69 difference in Elo. Hence, for the remainder of the external search experiments, we report results only for the mean scoring method. The full results are shown in Figure 10 in the appendix.
We then run a similar tournament including Ext-BoN: the results are shown in Table 2. The total number of games played in each respective tournament were 14689 for standard chess, 4480 for chess960, 2189 for Connect Four, and 6334 for Hex. In chess, MAV–MCTS with just 100 simulations achieves an internal Elo 68 higher than the base MAV model in chess, and generally performance rises consistently with more simulations, up to 340 internal Elo improvement at 2000 simulations. This is also illustrated in Figure 7(a), which demonstrates that Elo improves logarithmically as a function of the number of simulations.
In the case of Chess960, every game begins with a random position (among a preset 960 initial positions). We obtain comparable results to chess.
In the case of Connect Four, we notice that MAV–MCTS is particularly helpful in improving upon the base model. All external search agents improve by at least 244 Elo upon the base model. In Connect Four, we notice that improvements consistently rise with added simulations, albeit with smaller performance differences between simulation counts compared to chess. In Connect Four, we notice a relatively smaller performance gain between basic MCTS and the MAV model, and larger improvements between the MAV model and external search.
Since our final model was not adequately trained with state-tracking capabilities for Hex, in the Hex results only a game engine is used to determine terminal states, legal actions, and state transitions. In contrast to Connect Four, we notice a large improvement between basic MCTS and the MAV model in Hex, and improvements similar to chess from search. We suspect that this is due to the relative complexity of the games, but more research is needed to clarify the reason for these differences. It should be possible to support the engine-free logic in Hex similarly to chess going forward. Across all games, external MCTS performs consistently better as the number of simulations is increased.
@@ 5.4 Internal search results
For our internal search experiments, we vary the depth and breadth parameters of the minimax search, up to breadth 4 and depth 2. We map these different configurations into token counts by computing the average length of the prompt + response per configuration in our training data. Figure 7(b) shows the performance of the model as the function of token counts. We can see that playing strength, measured by external Elo [^2] For the internal search experiments, we didn’t play games until checkmate but stopped them once either side had reached a decisive, 1200+ centipawn advantage. This was done since the internal search training data did not train the model to predict the quickest win, which would sometimes result in the model aimlessly moving around pieces in winning positions., increases with the search budget.
@@ 6 Discussions and limitations
Despite promising results in the domain of perfect information board games indicating the potential of external and internal planning with LLMs, our initial study makes a number of assumptions that future work may need to address – namely, the ability to acquire or generate large quantities of game play data, as well as the availability of reliable solvers or game engines that can be used to annotate this data in order to create an appropriate training curriculum for the model. We have also yet to fully explore the distillation of more complex search procedures, which may introduce additional challenges.
@@
Meanwhile, we argue that there is potential utility in the separation of responsibilities within a larger system, and we demonstrate that it is possible to utilise the proposed models as tools within a larger generalist model, calling the specialist model through a natural language tool-use interface with an appropriately formatted input (objective, game dynamics, game state) and output (choice, justification). Specialist models may also be used as teachers for distilling their knowledge into generalists (Bounsi et al., 2024).
@@ Acknowledgements
Appendix A External MCTS: Additional Results
@@
Appendix B External MCTS: Algorithm Details

### v3

@@ 3.1 External Search with State-Tracking MAV
We adapt MCTS (see Algorithm 1) to use the learned MAV to predict the state transition from the string description of a parent node $\mathcal{N}_{t-1}(s)$ and action $a_{t-1}$ to the child node $\mathcal{N}_{t}$, as shown in Figure 3. We use the obtained information to store a (predicted) string description $s_{t}$ of the child node, which is then expanded with legal actions $\boldsymbol{a}_{t}^{L}$ and their associated state-action values $\boldsymbol{Q}^{(i)}(s_{t},\boldsymbol{a}_{t}^{L})$ used to compute a prior according to Equation 1 and the child’s value as the maximum over all state-action values, i.e., $\max_{a_{t}\in\boldsymbol{a}_{t}^{L}}Q^{(i)}(s_{t},a_{t})$.
@@
MAV may hallucinate and return responses that are improperly-formatted. To address this, for each game, a parser function is responsible for translating the output into the next state, legal actions, and their values. If the parsing fails – which can only happen if a response violates a predefined format – a special value of $\infty$- ∞ is assigned to the node to avoid future consideration. This procedure makes state-tracking MCTS robust to hallucinations, by explicitly avoiding states where the hallucination occurred.
3.2 Async MCTS and Dynamic Virtual Counts
Due to the heavy cost of inference in LLMs, we use an Asynchronous Monte Carlo tree search (Async MCTS) that queues multiple leaf evaluations simultaneously (i.e., in a “batch” of simulations). To avoid redundant concurrent computation, we employ a commonly-used optimization in parallel MCTS: virtual counts, which temporarily increase the visit counts by some constant $n_{c}$ during the simulation, which are then removed during the backpropagation phase. For more detail about this form of MCTS, please see Appendix A.
@@ 3.2 Async MCTS and Dynamic Virtual Counts
With low simulation counts, using the virtual losses accompanied by the fixed virtual counts $n_{c}$ did not strike a satisfactory balance between exploration and exploitation. To address this, we introduce the dynamic virtual counts that dynamically assign more weight to the virtual count values closer to the leaf nodes. Suppose simulation $m$ encounters leaf node $\mathcal{N}_{l}$. We define a virtual count for states-action pairs $(s_{t},a_{t})$ visited in simulation $m$ and leaf $\mathcal{N}_{l}$
@@ 4 Internal Search
The prompt for internal search resembles that of MAV, but includes a preamble with search parameters (tree depth and breadth, see Figure 4). The format of the target data is inspired by depth-first order traversal of minimax trees, as an iterative and linearized sequence of minimax tree traversal. Hence, following this format corresponds to an algorithmic execution task, akin to CLRS-Text (Markeeva et al., 2024).
The training data is based on target search trees, which were constructed using depth 3 (N.B., depth-zero is MAV), by annotating states (e.g., chess states being annotated with Stockfish), and expanding the top 5 moves into trees. This results in high quality target search trees, similar to those internally generated by game engines. To diversify training data and enable search budget control for trained models, prompts were composed with diverse search parameters, ranging depth 1–3 and breadth 2–5, and continuations yield corresponding trees that were subsampled to match the parameters while fitting into context size. This necessitated excluding the biggest parameter combination (i.e., omitting depth 3, breadth 5 examples).
An example internal search trace predicted by the trained internal search MAV (MAV–IS), along with the corresponding search tree, is shown in Figure 4.
@@
We leveraged the pre-trained MAV and fine-tuned it using a mixture of $60\%$ MAV data and $40\%$ search data. Fine-tuning was run for 20,000 steps, using a batch size of 512. Thus the model saw an order of magnitude more tokens during the MAV pre-training compared to the fine-tuning on internal search traces.
@@ 5.1 Evaluation
Evaluating LLMs in general and reasoning specifically is a broad field (Chang et al., 2024), within which games have been established as an evaluation benchmark (Costarelli et al., 2024; Duan et al., 2024). We evaluate MAV language models in a games league which head-to-head match-ups, sampling combinations uniformly at random from a pool of MAV and baseline agents. We report internal Elo ratings (relative Elo only between members of the population) as well as external Elo ratings where possible.
For chess evaluation, we rely on the state-of-the-art engine Stockfish at different playing strengths to estimate the widely used external Elo rating. This is done by calibrating the internal Elo with externally-reported ratings using a linear fit. As an additional, we include the Ext-BoN model (Ruoss et al., 2024a), which showed Grandmaster-level performance on chess with Transformers (see Appendix B for details). We use a set of Top Chess Engine Championship (TCEC) opening positions (tce, ) (see LABEL:tab:long in Appendix G). In each match-up between two agents, a specific opening is used, and agents swap seats to ensure each agent plays each opening both as black and as white. Every instance of Stockfish is run with 2 seconds of search time. We delineate further our evaluation setting together with the closest related works in Appendix B and Appendix C.
@@
Table 1: Performance comparison of different agents in Chess, Chess960, Connect Four, and Hex. In MAV–MCTS agents, $M$ refers to the number of simulations. The external Elos are estimated from the internal Elos achieved among agents. Important: The external Elos of the Stockfish agents are anchored to the CCLR Blitz Elos. Therefore, the Elo estimates for the MAV–MCTS agents are not accurate as those agents would not be able to play at blitz time controls.
@@
Table 2: Error rate analysis for MAV in chess positions. We consider positions drawn from Lichess puzzles, and from a dataset of randomly generated boards that never occur in the course of normal gameplay, and are highly out-of-distribution compared to the training data.
@@ 5.2 Multi-Action-Value Results
In terms of chess playing strength, MAV reaches an external Elo of 2923 when using mean scoring and 2875 when using max scoring. It thus outperforms several strong baselines, including Stockfish–L10 and performs comparably with the Ext-BoN model by Ruoss et al. (2024a). For Chess960 MAV outperforms Ruoss et al. (2024a), perhaps due to being trained explicitly on Chess960 data.
In Table 2, we analyze the legal move rate, the precision and recall of the predicted %top_all moves, and the accuracy of predicting the next FEN state. The results demonstrate that MAV is able to reliably perform all of these actions.
@@
During the opening and endgame, human players often rely on memorized opening lines and endgame theory, while middlegame requires more calculation and intuition. In Appendix D, we report the same pattern in the games of MAV and a concrete position where the model displays a creative play in unseen position. Overall, 10% of the positions played by MAV in evaluation appear in its training data, while between moves 20 and 50, virtually no position has been seen by MAV during training. These results show that in order to avoid losing games during middlegame, MAV is required to generalize.
@@ 5.3 External Search Results
Figure 5: Chess Elo improves with more test-time compute. Left: External search results, varying the number of simulations / MAV model calls. Right: Internal search results, varying the breadth ($b$) and depth ($d$) of the minimax search performed in a single model call. The $x$-axis shows the number of tokens in the prompt+response.
For our external-MCTS agents, we tune hyper-parameters using a combination of manual and head-to-head comparisons and report the details in Appendix A. We also include a basic MCTS baseline, which refers to MCTS with a uniform prior and random rollouts run with 100 simulations.
We first analyze the performance of the scoring method for the value function when used within MAV–MCTS. We do this by running a large tournament between various models with the details, including results shown in Figure 7 in Appendix A. MAV–MCTS with mean scoring performs noticeably better in all cases. Hence, for the remainder of the external search experiments, we report results only for the mean scoring method.
We then run an even bigger tournament including Ext-BoN with the results shown in Table 1. The total number of games played in each respective tournament were 14689 for Chess, 4480 for Chess960, 2189 for Connect Four, and 6334 for Hex. In chess, MAV–MCTS with just 100 simulations achieves an internal Elo 68 higher than the searchless MAV   and generally Elo performance improves logarithmically as a function of the number of simulations as illustrated in Figure 5 (right).
In the case of Chess960, where every game begins with a random position (among a preset 960 initial positions), we obtain comparable results to Chess. In the case of Connect Four, we notice that MAV–MCTS is particularly helpful in improving upon MAV, with all agents improving by at least 244 Elo. Similarly to Chess, the improvements consistently rise with added simulations, but here we notice a relatively smaller performance gain of MAV over a basic MCTS, and larger improvements of MAV–MCTS over MAV.
Since our final model was not adequately trained with state-tracking capabilities for Hex, in the Hex results only a game engine is used to determine legal actions, state transitions, and terminal states. In contrast to Connect Four, we notice a large improvement between basic MCTS and MAV in Hex, and improvements similar to chess from search. We suspect that this is due to the relative complexity of the games, but more research is needed to clarify the reason for these differences. It should be possible to support the engine-free logic in Hex similarly to Chess going forward. Across all games, external MCTS performs consistently better as the number of simulations is increased.
@@ 5.4 Internal Search Results
For our internal search experiments, we vary the depth and breadth parameters of the minimax search, up to breadth 4 and depth 2. We map these different configurations into token counts by computing the average length of the prompt + response per configuration in our training data. Figure 5 (right) shows the performance of MAV–IS as the function of token counts. We can see that playing strength, measured by external Elo [^3] For the internal search experiments, we stopped games once either side had reached a decisive, 1200+ centipawn advantage. This was done since the internal search training data did not include quickest wins, which would sometimes result in the model aimlessly moving pieces around in winning positions., increases with the search budget.
We also analyzed if the MAV–IS search produces the tree of the requested shape whenever possible, considering the available number of legal moves at each node. MAV–IS constructed an accurately shaped tree 99.6% of the time on a sample 5980 positions taken from existing games.
@@ 6 Discussions and Limitations
Despite promising results in the domain of perfect information board games indicating the potential of external and internal planning with LLMs, our initial study makes a number of assumptions that future work may need to address – namely, the ability to acquire or generate large quantities of game play data, as well as the availability of reliable solvers or game engines that can be used to annotate this data in order to create an appropriate training curriculum for the model.
@@ Acknowledgements
Appendix A External MCTS
A.1 Additional Plots and Results
(a) An overview of Async MCTS.
(b) Dynamic virtual counts for $n_{\text{max}}=8$ and $n_{\text{min}}=2$. Virtual counts from separate simulations are added, as is the case in the $+4$ node with two children ($+4=(+2)+(+2)$).
Figure 6: Async MCTS and Dynamic Virtual Counts building blocks.
@@ A.1 Additional Plots and Results
A.2 Performance Considerations and Async MCTS
Since modern large transformer models can take orders of magnitude longer for a single inference than an optimized deep convolutional network (such as the ones used in MuZero (Schrittwieser et al., 2020)), keeping the elapsed time for an MCTS search to a reasonable level can be a challenge. The inference time of a single model call dominates the time take for a search, so we aim to keep the total calls low. Since the prior and value for a new node are both obtained for the MAV, there is at most one model call per simulation of MCTS. In the low simulation count regime, reusing the search trees is a simple optimization that reduces the number of server calls by 5-10% in our experiments.
Despite this lower number of inference calls, it can still be costly to make many MAV calls. Hence, we run an asynchronous version of MCTS that is conceptually very similar to Algorithm 1, but with the main loop changed. An overview is shown in Figure 6.
Parallel MCTS implementations usually rely on virtual losses (Mirsoleimani et al., 2017; Chaslot et al., 2008) to avoid multiple threads choosing the same simulation paths down due to deterministic action choices. Our Async MCTS introduces three new parameters:
Batch size $b$.
The number of simulations, each leading to an evaluation (MAV call), that are done asynchronously at any given time.
Timeout $t_{0}$.
Virtual Count Value $n_{c}$.
The “virtual” value to temporarily add to a visit count $N(s,a)$ during the simulation to discourage all simulations in a batch from choosing the same path.
In Async MCTS, simulations are instead performed in batches. First, $b$ downward passes are performed serially to queue $b$ evaluations. The $b$ evaluations are performed in parallel and the MAV outputs are returned asynchronously.
The evaluations are then processed one-by-one as they return. For each queued evaluation that was successful (i.e.,  did not time out), its values are then backpropagated up the tree (serially). The state-action pairs that are visited before the queuing of evaluations are assigned temporary virtual counts: their visit counts are temporary set to $N(s,a)=N(s,a)+n_{c}$. The visit counts and virtual counts are then subtracted upon the upward pass, i.e., $N(s,a)=N(s,a)-n_{c}$.
The average amount of time taken per search is shown in Table 1.
Number of Simulations, $M$ Average Elapsed Time per Search (seconds)
100 39.82
250 114.12
500 187.10
1000 378.32
2000 729.23
Table 3: Average elapsed time taken per search by Async MCTS.
A.3 Hyper-Parameter Tuning and Values
For our external-MCTS agents, we tune hyper-parameters using a combination of manual and head-to-head comparisons. We found that setting $\tau=0.1$, $k=5$, and $\varepsilon=0.05$ worked best for the prior (Equation 1).
For Async MCTS we used a batch size $b=16$. We found that lower batch sizes were always preferred in terms of performance. We used a timeout of $t_{0}=60$ seconds, and (only as a baseline) static virtual loss parameter $n_{c}=10$. In head-to-head comparisons, we found that a tuned dynamic virtual count would generally outperform static virtual counts. We use the robust child strategy for final move selection Chaslot et al. (2009) (choosing the action with the highest number of visits), using the total reward as a tie-breaker.
We used dynamic virtual count values $n_{\text{min}}=2$ in all cases, and $n_{\text{max}}$ varying with the number of simulations: $(M,n_{\text{max}})\in$$(100,8)$, $(250,8)$, $(500,8)$, $(1000,16)$, (2000,32)}(2000,32)\}( 2000 , 32 ) }.
A.4 External MCTS: Algorithm Details